An investment project provides cash inflows of $620 per year for 11 years.

A. What is the project payback period if the initial cost is 1,240?

B. What is the project payback period if the initial cost is 4,216?

C.

What is the project payback period if the initial cost is 7,440?
Answer #1

Payback period=initial investment/annual cash flows

1.Payback period=(1240/620)=2 years

2.Payback period=(4216/620)=6.8 years

3.Payback period=(7440/620)=0(no payback).